OpenStudy (anonymous):

Solve the problem. Find all values of k so that the given points are\sqrt {29} units apart. (-5, 5), (k, 0) Choose the right answer a. 3, 7 b. 7 c. -3, -7 d. -7

OpenStudy (anonymous):

\[D = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\]

OpenStudy (anonymous):

(k+5)^2 + (0-5)^2 = 29 k^2 +10k+25 +25 -29=0 k^2+10k+21 = 0 k = -3 or -7

OpenStudy (anonymous):

so answer is c..........is it still a bad answer?

OpenStudy (anonymous):

I think you are right... saruz, thanks i will see.

OpenStudy (anonymous):

¿Por qué no haces clic en una buena respuesta, entonces?

OpenStudy (anonymous):

Tu hablas español?

OpenStudy (anonymous):

No, no hablo español. Hay un pequeño truco ... estoy usando el software ... Yo no soy de España

OpenStudy (anonymous):

Ok, can you keep helping me some minutes more?

OpenStudy (anonymous):

i cant say i can help but i will definitely try my best

OpenStudy (anonymous):

Find the value of the indicated trigonometric function of the angle θ in the figure. Give an exact answer with a rational denominator.